1. Derive and solve the equation of motion of a particle, in a uniform gravitational field, projected with initial horizontal velocity v 0 at a height h.
  2. Write the Lagrangian of this particle. Show that the Euler-Lagrange equations of motion for this particle is identical to what one would obtain from Newton’s second law.
1
Expert's answer
2021-08-16T08:37:05-0400

The Newton's second law says

"m{\bf a}={\bf F}"

"m{\ddot x}=0,\\
m{\ddot y}=-mg."

Equations of motion:

"{\ddot x}=0,\quad {\dot x}(0)=v_0,\quad x(0)=0,\\{\ddot y}=-g,\quad {\dot y}(0)=0,\quad y(0)=h."

Solutions:

"x(t)=v_0t,\\
y(t)=h-\frac{gt^2}{2}"

2. The lagrangian of the system

"L=T-V=\frac{m({\dot x}^2+{\dot y}^2)}{2}-mgy"

The Euler-Lagrange equations

"\frac{d}{dt}\left(\frac{\partial L}{\partial {\dot x}}\right)-\frac{\partial L}{\partial {x}}=0,\\\frac{d}{dt}\left(\frac{\partial L}{\partial {\dot y}}\right)-\frac{\partial L}{\partial {y}}=0,"

give

"m{\ddot x}=0,\\
m{\ddot y}=-mg."